# QVGen: Pushing the Limit of Quantized Video Generative Models

### v1

@@ 4.2 Progressively Shrinking $\Phi$ via Rank-Decay
To improve QAT while eliminating the inference overhead, we propose to progressively remove $\Phi$ throughout the training process. This allows the model to benefit from $\Phi$ during QAT, while ultimately yielding a standard quantized model [35, 8] with no extra inference cost. To achieve this goal, a straightforward solution is to decay all parameters of $\Phi$ directly. However, we have noticed that it is ineffective and suboptimal (see Tab. 5.3). This calls for a fine-grained decay strategy.
(a) CogVideoX-2B [72]
(b) Wan 1.3B [64]
Figure 4: Singular value variation in ${\mathbf{W}}_{\Phi}$ across training iterations for 4-bit video DMs. We visualize the average of the singular values $\{\sigma_{s}\}_{s=1,2,\ldots,2^{10}}\cup\{\sigma_{d}\}$ across layers of all Attention blocks [62] and feed-forward networks (FFNs), respectively. “00 step” denotes the initialization state before the start of QAT.
Based on the above analysis, we investigate the contribution of fine-grained components in $\Phi$. Specifically, we apply singular value decomposition (SVD) [32, 36] to ${\mathbf{W}}_{\Phi}$ at various training steps:
@@
where $d=\min\{n,m\}$ and $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{d}$ are the singular values. The vectors ${\bm{u}}_{s}$ and ${\bm{v}}_{s}$ denote the left and right singular vectors associated with $\sigma_{s}$, respectively. By tracking the evolution of the average $\sigma_{s}$, we observe two key findings (exemplified by Fig. 4 (a) Attention):
[…]
These findings suggest that an increasing number of orthonormal directions $\{{\bm{u}}_{s},{\bm{v}}_{s}\}$ contribute little, as their associated singular values $\sigma_{s}$ are small [75, 71]. Hence, as training proceeds, only an increasingly low-rank portion of $\Phi$ (i.e., $\Sigma^{d^{\prime}}_{s=1}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{T}$, where $d^{\prime}<d$) is needed, and the remaining components can be decayed without noticeably affecting performance. Motivated by this, we propose a novel rank-decay schedule that progressively shrinks $\Phi$ by repeatedly identifying and eliminating the above-mentioned low-impact parts. First, to attain these parts, we reformulate the computation of $\Phi$ as:
@@
Here, $u$ follows a cosine annealing schedule that decays from $1$ to 00, $\lambda\in(0,1]$ represents the shrinking ratio, and $\odot$ denotes element-wise multiplication. Eq. (13) and Eq. (14) allow us to progressively eliminate the low-impact components of $\Phi$ (i.e., $[\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{u}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]$ and $[\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{v}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{T}$). Once $u$ reaches 00, we truncate $\{{\mathbf{L}},{\mathbf{R}}\}$ to $\{{\mathbf{L}}^{\prime},{\mathbf{R}}^{\prime}\}$, and rewrite ${\mathbf{W}}_{\Phi}$ as:
@@
In Eq. (15), the rank of ${\mathbf{W}}_{\Phi}$ is shrunk from $r$ to $(1-\lambda)r$. During the subsequent training phase, the above procedures (i.e., both decomposition and decay) are iteratively applied. Ultimately, we fully eliminate $\Phi$ by reducing $r$ to 00, which incurs negligible impact on model performance (see Tab. 5.3). It is worth noting that we set $\lambda=\frac{1}{2}$ for Eq. (14) in this work, based on its effectiveness demonstrated in Tab. 5.3. Overall, the overview of the rank-decay schedule is exhibited in Fig. 2 (b).
@@ 5.1 Implementation Details
Models. We conduct experiments on open-source SOTA video DMs, including CogVideoX-2B and 1.5-5B [72], and Wan 1.3B and 14B [64]. Classifier-free guidance (CFG[21] is used for all models, and the frame number of generated videos is fixed to $49$ for CogVideoX-2B and $81$ for the others.
Baselines. We adopt previous powerful PTQ and QAT methods as baselines: ViDiT-Q [76], SVDQuant [32], LSQ [8], Q-DM [35], and EfficientDM [17]. Since these methods were designed for image DMs or convolutional neural networks (CNNs), we adapt these works to video DMs using their open-source code (if available) or the implementation details provided in the corresponding papers. Without specific clarification, static per-channel weight quantization with dynamic per-token activation quantization, a common practice in the community [76, 41], is used for all the linear layers.
Training. We employ $16K$ captioned videos from OpenVidHQ-4M [48] as the training dataset. The AdamW [42] optimizer is utilized with a weight decay of $10^{-4}$. We employ a cosine annealing schedule to adjust the learning rate over training. During QAT, we train Wan 14B [64] and CogVideoX1.5-5B [72] for $16$ epochs on $10032\times H$ 100 GPUs and $10016\times H$ 100 GPUs, respectively. For the other DMs, we employ $8$ training epochs on $1008\times H$ 100 GPUs. Additionally, we allocate the same training iterations for each decay phase (i.e., shrinking the remaining $r$ to $(1-\lambda)r$). The same training settings are applied to all QAT baselines.
Evaluation. We select $8$ dimensions in VBench [25] with unaugmented prompts to comprehensively evaluate the performance following previous studies [76, 52]. Moreover, for quantized huge ($\geq$ 5B parameters) DMs, we additionally report the results on VBench-2.0 [78] with augmented prompts to measure the adherence of videos to physical laws, commonsense reasoning, etc.
More detailed experimental setups can be found in the appendix.
Table 1: Performance comparison across different quantization methods on VBench [25]. “†” indicates PTQ methods and “*” signifies QAT methods. “Full Prec.” denotes the BF16 model. $\clubsuit$” represents that we apply fine-grained per-group weight-activation quantization with a group size of 64 and keep some linear layers unquantized, which is the same as the official settings of SVDQuant [32] (details can be found in the appendix). Best and second-best results are highlighted in bold and underline formats, respectively. The gaps between them are marked in red subscripts.
{tabu}
[t!]l|c|llllllll Method #Bits (W/A) Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
CogVideoX-2B ($\texttt{CFG}=6.0,480p,\texttt{fps}=8$)
@@
ViDiT-Q [76] 4/6 54.72 43.01 92.18 43.22 90.76 81.02 26.25 20.41
SVDQuant [32] 4/6 58.27 47.06 95.28 40.83 92.41 87.45 27.69 21.34
SVDQuant [32]†♣ 4/4 51.60 49.40 97.69 42.22 94.03 91.78 25.67 22.89
LSQ [9] 4/4 58.73 54.20 97.57 45.00 92.97 92.41 24.06 23.17
Q-DM [35] 4/4 54.96 52.71 98.00 48.61 93.82 91.86 28.02 23.87
EfficientDM [17] 4/4 55.96 51.97 98.03 46.67 94.10 91.70 27.76 24.28
QVGen (Ours) 4/4 60.16+1.43 54.61+0.41 98.06+0.03 67.22+18.61 94.38+0.28 93.01+0.60 31.42+3.40 24.61+0.33
LSQ [9] 3/3 56.46 40.35 97.98 0.56 94.08 89.18 4.80 13.80
Q-DM [35] 3/3 50.88 40.41 98.03 5.56 93.93 87.75 7.33 15.98
EfficientDM [17] 3/3 52.86 44.58 97.13 28.61 93.15 88.26 15.42 20.42
QVGen (Ours) 3/3 58.36+1.90 50.54+5.96 98.37+0.34 53.89+25.28 94.55+0.47 90.50+1.32 23.85+8.43 22.92+2.50
Wan 1.3B ($\texttt{CFG}=5.0,480p,\texttt{fps}=16$)
@@
ViDiT-Q [76] 4/6 56.24 50.18 94.81 52.43 89.67 82.53 13.45 19.58
SVDQuant [32] 4/6 58.16 51.27 97.05 49.44 93.74 91.71 14.18 23.26
SVDQuant [32]†♣ 4/4 57.57 46.30 94.21 72.22 93.16 77.96 12.73 21.91
LSQ [9] 4/4 59.11 49.09 98.35 71.11 92.66 91.67 10.38 18.83
Q-DM [35] 4/4 60.40 52.50 97.22 76.67 93.37 89.26 13.28 21.63
EfficientDM [17] 4/4 60.70 53.57 96.18 56.39 93.74 91.70 11.77 21.19
QVGen (Ours) 4/4 63.08+2.38 54.67+1.10 98.25-0.10 77.78+1.11 94.08+0.34 92.57+0.87 15.32+2.04 23.01+1.38
LSQ [9] 3/3 58.80 46.86 98.22 23.61 91.86 89.42 0.89 15.51
Q-DM [35] 3/3 56.19 44.95 95.13 76.94 92.09 83.82 1.79 16.89
EfficientDM [17] 3/3 42.32 33.52 96.50 70.28 92.10 74.79 0.04 11.38
QVGen (Ours) 3/3 67.35+8.55 49.71+2.85 98.93+0.71 84.14+7.20 93.62+1.52 92.25+2.83 5.71+3.92 20.11+3.22
@@ 5.2 Performance Analysis
Comparison with baselines. We report VBench [25] score comparisons in Tab. 5.1. In W4A4 quantization, recent QAT methods [8, 17, 35] show non-negligible performance degradation. With W3A3, performance drops become more pronounced. In contrast, the proposed QVGen achieves substantial performance recovery in 3-bit models and comparable results to full-precision models in 4-bit quantization. Specifically, it shows higher scores or less than a $2\%$ decrease in all metrics for W4A4 CogVideoX-2B [72], except Scene Consistency. For PTQ baselines [76, 32], all fail to generate meaningful content in W4A4 per-channel and per-token settings. Therefore, we apply W4A6 quantization for these methods and also utilize fine-grained per-group W4A4 quantization for SVDQuant [32]. In these cases, W4A4 QVGen outperforms them by a large margin, particularly with $8.37$ and $14.61$ higher Aesthetic Quality and Subject Consistency for Wan 1.3B compared to W4A4 SVDQuant. In addition to these findings, we observe that for Wan 1.3B, Dynamic Degree recovers easily during QAT, even surpassing that of the full-precision model. However, for CogVideoX-2B, this metric significantly drops. Moreover, Scene Consistency is the most challenging metric to maintain across models and methods.
Beyond the quantitative results, we provide qualitative results in Fig. 1 and the appendix, where QVGen markedly improves visual quality over prior methods. Although a clear gap remains between 3-bit and 4-bit outputs, our work lays a foundation for practical 3-bit or lower video DM quantization.
Table 2: Performance for huge video DMs on VBench [25].
{tabu}
[t!]l|c|llllllll Method #Bits (W/A) Imaging Quality $\uparrow$ Aesthetic Quality $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Background Consistency $\uparrow$ Subject Consistency $\uparrow$ Scene Consistency $\uparrow$ Overall Consistency $\uparrow$
CogVideoX1.5-5B ($\texttt{CFG}=6.0,720p,\texttt{fps}=16$)
@@
(c) CogVideoX1.5-5B [72]
(d) Wan 14B [64]
Figure 5: Performance for huge video DMs on VBench-2.0 [78]. Our 4-bit models exhibit a minimal drop of $\sim 1\%$ in total score.
Results for huge DMs. To demonstrate the scalability of our method, we further test two huge video DMs, including CogVideoX1.5-5B and Wan 14B, at $720p$ resolution. As illustrated in Tab. 5.2, our 3-bit and 4-bit models follow the same pattern seen in smaller models (Tab. 5.1). However, 3-bit quantization incurs much larger drops on demanding metrics such as Scene and Overall Consistency, underscoring the challenge of pushing these larger models to 3 bits. In Fig. 5, we further assess the models with VBench-2.0 [78]; the W4A4 DMs incur only negligible overall performance loss.

### v3

@@ 3.2 Progressively Shrinking $\Phi$ via Rank-Decay
To improve QAT while eliminating the inference overhead, we propose to progressively remove $\Phi$ throughout the training process. This allows the model to benefit from $\Phi$ during QAT, while ultimately yielding a standard quantized model (Li et al., 2023b; Esser et al., 2020b) with no extra inference cost. To achieve this goal, a straightforward solution is to decay all parameters of $\Phi$ directly. However, we have noticed that it is ineffective and suboptimal (see Tab. 5). This calls for a fine-grained decay strategy.
(a) CogVideoX-22(Yang et al., 2025)
(b) Wan 1.31.3(WanTeam et al., 2025)
Figure 4: Singular value variation in ${\mathbf{W}}_{\Phi}$ across training iterations for $4$-bit video DMs. We visualize the average of the singular values $\{\sigma_{s}\}_{s=1,2,\ldots,2^{10}}\cup\{\sigma_{d}\}$ across layers of all Attention blocks (Vaswani et al., 2023) and feed-forward networks (FFNs), respectively. “0 step” denotes the initialization state before QAT.
Based on the above analysis, we investigate the contribution of fine-grained components in $\Phi$. Specifically, we apply singular value decomposition (SVD) (Li et al., 2025; 2023c) to ${\mathbf{W}}_{\Phi}$ at various training steps:
@@
where $d=\min\{n,m\}$ and $\sigma_{1}\geq\sigma_{2}\geq\ldots\geq\sigma_{d}$ are the singular values. The vectors ${\bm{u}}_{s}$ and ${\bm{v}}_{s}$ denote the left and right singular vectors associated with $\sigma_{s}$, respectively. By tracking the evolution of the average $\sigma_{s}$, we observe two key findings (exemplified by Fig. 4 (a) Attention): (i) ${\mathbf{W}}_{\Phi}$ contains a substantial number of small singular values. For example, approximately $73\%$ (0-th step) of the average $\sigma_{s}$ are $\sim 14\times$ smaller than the largest one $\sigma_{1}$; (ii) The presence of these small $\sigma_{s}$ becomes increasingly pronounced as QAT progresses, with the proportion rising to $99\%$ ($2K$-th step).
These findings suggest that an increasing number of orthonormal directions $\{{\bm{u}}_{s},{\bm{v}}_{s}\}$ contribute little, as their associated singular values $\sigma_{s}$ are small (Zhang et al., 2015; Yang et al., 2020). Hence, as training proceeds, only an increasingly low-rank portion of $\Phi$ (i.e., $\Sigma^{d^{\prime}}_{s=1}\sigma_{s}{\bm{u}}_{s}{\bm{v}}_{s}^{T}$, where $d^{\prime}<d$) is needed, and the remaining components can be decayed without noticeably affecting performance. Motivated by this, we propose a novel rank-decay schedule that progressively shrinks $\Phi$ by repeatedly identifying and eliminating the above-mentioned low-impact parts. First, to attain these parts, we reformulate the computation of $\Phi$ as:
@@
Here, $u$ follows a cosine annealing schedule that decays from $1$ to 0, $\lambda\in(0,1]$ represents the shrinking ratio, and $\odot$ denotes element-wise multiplication. Eq. (12) and Eq. (13) allow us to progressively eliminate the low-impact components of $\Phi$ (i.e., $\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{u}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{u}}_{r}]$ and $\sqrt{\sigma_{(1-\lambda)r+1}}{\bm{v}}_{(1-\lambda)r+1},\ldots,\sqrt{\sigma_{r}}{\bm{v}}_{r}]^{T}$). Once $u$ reaches 0, we truncate $\{{\mathbf{L}},{\mathbf{R}}\}$ to $\{{\mathbf{L}}^{\prime},{\mathbf{R}}^{\prime}\}$, and rewrite ${\mathbf{W}}_{\Phi}$ as:
@@
In Eq. (14), the rank of ${\mathbf{W}}_{\Phi}$ is shrunk from $r$ to $(1-\lambda)r$. During the subsequent training phase, the above procedures (i.e., both decomposition and decay) are iteratively applied. Ultimately, we fully eliminate $\Phi$ by reducing $r$ to 0, which incurs negligible impact on model performance (see Tab. 3). It is worth noting that we set $\lambda=\frac{1}{2}$ for Eq. (13) in this work, based on its effectiveness demonstrated in Tab. 5. Overall, the overview of the rank-decay schedule is exhibited in Fig. 2 (b).
@@ 4.1 Implementation Details
Models. We conduct experiments on open-source SOTA video DMs, including CogVideoX-22B and 1.5-55(Yang et al., 2025), and Wan 1.31.3B and 1414(WanTeam et al., 2025). Classifier-free guidance (CFG(Ho & Salimans, 2022) is used for all models, and the frame number of generated videos is fixed to $49$ for CogVideoX-22B and $81$ for the others.
Baselines. We adopt previous powerful PTQ and QAT methods as baselines: ViDiT-Q (Zhao et al., 2025a), SVDQuant (Li et al., 2025), LSQ (Esser et al., 2020b), Q-DM (Li et al., 2023b), and EfficientDM (He et al., 2024). Since these methods were designed for image DMs or convolutional neural networks (CNNs), we adapt these works to video DMs using their open-source code (if available) or the implementation details provided in the corresponding papers. Without specific clarification, static per-channel weight quantization with dynamic per-token activation quantization, a common practice in the community (Zhao et al., 2025a; Liu et al., 2023), is used for all linear layers.
Training. We employ $16K$ captioned videos from OpenVidHQ-4M (Nan et al., 2025) as the training dataset. The AdamW (Loshchilov & Hutter, 2019) optimizer is utilized with a weight decay of $10^{-4}$. We employ a cosine annealing schedule to adjust the learning rate over training. During QAT, we train Wan 1414(WanTeam et al., 2025) and CogVideoX1.5-55(Yang et al., 2025) for $16$ epochs on $32\times H100$ GPUs and $16\times H100$ GPUs, respectively. For the other DMs, we employ $8$ training epochs on $8\times H100$ GPUs. Additionally, we allocate the same training iterations for each decay phase (i.e., shrinking the remaining $r$ to $(1-\lambda)r$). The same settings are applied to all QAT baselines.
Evaluation. We select $8$ dimensions in VBench (Huang et al., 2024b) with unaugmented prompts to comprehensively evaluate the performance following previous studies (Zhao et al., 2025a; Ren et al., 2024). Moreover, for huge ($\geq$ 55B parameters) DMs, we additionally report the results on VBench-2.0 (Zheng et al., 2025a) with augmented prompts to measure the adherence of videos to physical laws, reasoning, etc. More detailed experimental setups can be found in Sec. D.
Table 1: Performance comparison across different quantization methods on VBench (Huang et al., 2024b). “†” indicates PTQ methods and “*” signifies QAT methods. “Full Prec.” denotes the BF16 model. $\clubsuit$” represents that we apply fine-grained per-group weight-activation quantization with a group size of $64$ and keep some linear layers unquantized, which is the same as the official settings of SVDQuant (Li et al., 2025) (details can be found in Sec. D). Best and second-best results are highlighted in bold and underline formats, respectively.
Method
#Bits
(W/A)
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
CogVideoX-22B ($\texttt{CFG}=6.0,480p,\texttt{fps}=8$)
@@
ViDiT-Q (Zhao et al., 2025a) 4/6 54.72 43.01 92.18 43.22 90.76 81.02 26.25 20.41
SVDQuant (Li et al., 2025) 4/6 58.27 47.06 95.28 40.83 92.41 87.45 27.69 21.34
SVDQuant (Li et al., 2025)†♣ 4/4 51.60 49.40 97.69 42.22 94.03 91.78 25.67 22.89
LSQ (Esser et al., 2020a) 4/4 58.73 54.20 97.57 45.00 92.97 92.41 24.06 23.17
Q-DM (Li et al., 2023b) 4/4 54.96 52.71 98.00 48.61 93.82 91.86 28.02 23.87
EfficientDM (He et al., 2024) 4/4 55.96 51.97 98.03 46.67 94.10 91.70 27.76 24.28
QVGen (Ours) 4/4 60.16 54.61 98.06 67.22 94.38 93.01 31.42 24.61
LSQ (Esser et al., 2020a) 3/3 56.46 40.35 97.98 0.56 94.08 89.18 4.80 13.80
Q-DM (Li et al., 2023b) 3/3 50.88 40.41 98.03 5.56 93.93 87.75 7.33 15.98
EfficientDM (He et al., 2024) 3/3 52.86 44.58 97.13 28.61 93.15 88.26 15.42 20.42
QVGen (Ours) 3/3 58.36 50.54 98.37 53.89 94.55 90.50 23.85 22.92
Wan 1.31.3B ($\texttt{CFG}=5.0,480p,\texttt{fps}=16$)
@@
ViDiT-Q (Zhao et al., 2025a) 4/6 56.24 50.18 94.81 52.43 89.67 82.53 13.45 19.58
SVDQuant (Li et al., 2025) 4/6 58.16 51.27 97.05 49.44 93.74 91.71 14.18 23.26
SVDQuant (Li et al., 2025)†♣ 4/4 57.57 46.30 94.21 72.22 93.16 77.96 12.73 21.91
LSQ (Esser et al., 2020a) 4/4 59.11 49.09 98.35 71.11 92.66 91.67 10.38 18.83
Q-DM (Li et al., 2023b) 4/4 60.40 52.50 97.22 76.67 93.37 89.26 13.28 21.63
EfficientDM (He et al., 2024) 4/4 60.70 53.57 96.18 56.39 93.74 91.70 11.77 21.19
QVGen (Ours) 4/4 63.08 54.67 98.25 77.78 94.08 92.57 15.32 23.01
LSQ (Esser et al., 2020a) 3/3 58.80 46.86 98.22 23.61 91.86 89.42 0.89 15.51
Q-DM (Li et al., 2023b) 3/3 56.19 44.95 95.13 76.94 92.09 83.82 1.79 16.89
EfficientDM (He et al., 2024) 3/3 42.32 33.52 96.50 70.28 92.10 74.79 0.04 11.38
QVGen (Ours) 3/3 67.35 49.71 98.93 84.14 93.62 92.25 5.71 20.11
@@ 4.2 Performance Analysis
Comparison with baselines. We report VBench score comparisons in Tab. 1. In W4A4 quantization, recent QAT methods (Esser et al., 2020b; He et al., 2024; Li et al., 2023b) show non-negligible performance degradation. With W3A3, performance drops become more pronounced. In contrast, the proposed QVGen achieves substantial performance recovery in $3$-bit models and comparable results to full-precision models in $4$-bit quantization. Specifically, it shows higher scores or less than a $2\%$ decrease in all metrics for W4A4 CogVideoX-22(Yang et al., 2025), except Scene Consistency. For PTQ baselines (Zhao et al., 2025a; Li et al., 2025), all fail to generate meaningful content in W4A4 per-channel and per-token settings. Therefore, we apply W4A6 quantization for these methods and also utilize fine-grained per-group W4A4 quantization for SVDQuant (Li et al., 2025). In these cases, W4A4 QVGen outperforms them by a large margin, particularly with $8.37$ and $14.61$ higher Aesthetic Quality and Subject Consistency for Wan 1.31.3B compared to W4A4 SVDQuant. In addition to these findings, we observe that for Wan 1.31.3B, Dynamic Degree recovers easily during QAT, even surpassing that of the BF16 model. However, for CogVideoX-22B, this metric significantly drops. Moreover, Scene Consistency is the most challenging metric to maintain across models and methods.
Beyond the quantitative results, we provide qualitative results in Fig. 1 and Sec. L, where QVGen markedly improves visual quality over prior methods. Although a clear gap remains between $3$-bit and $4$-bit outputs, our approach shifts the Pareto frontier toward higher accuracy at a smaller model size than existing techniques (e.g., $3$-bit QVGen for CogVideoX achieves a superior Aesthetic Quality than previous $4$-bit methods in Tab. 1). We view this initial exploration as setting a direction and providing valuable insight for future work on 22‑bit quantization.
Besides, we provide more comparisons with additional metrics in Sec. H and comparisons under relatively higher bit-width in Sec. E. We also apply our approach to image generation in Sec. K.
Table 2: Performance for huge video DMs on VBench. Comparison with baselines can be found in Sec. I.
Method
#Bits
(W/A)
Imaging
Quality
$\uparrow$
Aesthetic
Quality
$\uparrow$
Motion
Smoothness
$\uparrow$
Dynamic
Degree
$\uparrow$
Background
Consistency
$\uparrow$
Subject
Consistency
$\uparrow$
Scene
Consistency
$\uparrow$
Overall
Consistency
$\uparrow$
CogVideoX1.5-55B ($\texttt{CFG}=6.0,720p,\texttt{fps}=16$)
@@
Results for huge DMs. To demonstrate the scalability of our method, we further test two huge video DMs, including CogVideoX1.5-55B and Wan 1414B, at $720p$ resolution. As illustrated in Tab. 2, our $3$-bit and $4$-bit models follow the same pattern seen in smaller models (Tab. 1). However, $3$-bit quantization incurs much larger drops on demanding metrics such as Scene and Overall Consistency,
(a) CogVideoX1.5-55(Yang et al., 2025)
(b) Wan 1414(WanTeam et al., 2025)
Figure 5: Performance for huge video DMs on VBench-2.0 (Zheng et al., 2025a). Our 44-bit models exhibit a minimal drop of $\sim 1\%$ in total score.
underscoring the challenge of pushing these larger models to $3$ bits. In Fig. 5, we further assess the models with VBench-2.0; the W4A4 DMs incur only negligible overall performance loss.